OpenStudy (anonymous):

(1/3)[6x - 2(x - 4)] = 8

OpenStudy (lgbasallote):

multiply both sides by 3 first...what do you get?

OpenStudy (anonymous):

{6x-2(x-4)}=24

OpenStudy (lgbasallote):

right. now distribute -2 into (x-4)

OpenStudy (anonymous):

{6x-2x+8}=24 {4x+8}=24 4x=16 x=4

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

so there's your answer

OpenStudy (anonymous):

Thanks again :)